# SPAND: Sleep Prediction Architecture using Network Dynamics

### 20 Jan 2024 v1

@@ 4.1. Problem description
We pose a problem of predicting sleep labels for the next day based on wearable and mobile data for the current and past few days. Formally, the multi-modal data with $m$ features for a given day $k$ is represented by the vector $x[k]\in\mathbb{R}^{m}$ and the corresponding sleep duration in minutes is represented by $s[k]$. The feature matrix $X[k]$ is composed of past $L$ days of feature data, where $L$ is an indicator of temporal memory, and the sleep label $y[k]$ is generated from sleep duration $s[k]$ (mins). The recommended sleep duration for young adults is between $8-10$ hours(Paruthi et al., 2016)(Hirshkowitz et al., 2015b)(Hirshkowitz et al., 2015a), so we choose a threshold of 8 hours for differentiating between good and poor sleep behavior. Also, the sleep duration distribution in the dataset is relatively balanced around 8 hours. The sleep duration for our collected dataset had a mean of 7.2 hours with a standard deviation of 2.1 hours.
@@ 4.2. Graph Network Construction
We validated social interactions among study participants using data from pre- and post-study social network surveys. These data allowed us to quantify the closeness between participants and create graph networks, which are detailed in Supplementary Material section D.2. Our findings affirm the presence of social interactions and phone communication among the study participants.
@@
4.3. SPAND: Proposed Architecture
@@ 4.3.2. Attention-based Context Aggregation
To integrate graph structure in this attention mechanism, these coefficients are only computed for the first-hop neighbors. In addition, to compare different neighbors and have stable information aggregation, the coefficients are normalized across all neighbors. Finally, the normalized coefficients are used to compute a weighted aggregation of features from all over the node $u_{i}$’s neighbors,
@@ 4.3.5. End-to-End Model
Figure 3. SPAND Architecture. The top module based on GAN takes both the feature matrix and graph as input. The layer operates on graphs by applying attention shown by blue dashed links and original graph links are shown in solid green lines. The bottom module takes sequential data as input and extracts temporal trends. Finally, embedded features from both modules are combined in the aggregation module that produces predicted sleep labels.
The complete architecture of SPAND is shown in Fig.3. The input to the model is a 3-tuple object $I=(X[k],S^{L}_{k},A_{N}^{k})$, where $X[k]=\left[x_{1}[k],x_{2}[k],...x_{N}[k]\right]$ is a matrix containing feature vectors of $N$ participants each represented by $x_{i}\in\mathbb{R}^{m}$ for day indexed by $k$. The sequence data fed to the LSTM unit is represented by $S^{L}_{k}$. The graph network corresponding to the $N$ participants in feature matrix $X[k]$ is represented by adjacency matrix $A_{N}^{k}$. This corresponding data label is sleep duration for all $N$ participants for the next day represented by $y[k+1]$. After extracting temporal dynamics and multi-participant aggregated features in an embedded space, the output of both modules is normalized and concatenated before being fed to a small network of dense layers that produce the final output.
@@ 4.4.2. Machine Learning Pipeline
  • Random Split: The dataset is randomly split into training and test sets. We use random splitting to avoid selection bias. Around $50\%$ of data is used for training, $10\%$ for validation, and the remaining $40\%$ for testing with SMS graphs. The same training data is used to train all 4 models and then the same test data is used for testing across all models. This ensures that a fair comparison is conducted between the models. In order to mitigate the impact of model sensitivity to parameter initialization and depict the stability of the model, we use bootstrapping(Efron and Tibshirani, 1994)(Efron, 1992) and repeat the training and testing procedure 25 times and report average results along with the standard deviation across trials.
@@
To reduce the computation time, we use early stopping in the training process with a patience value of 30 epochs and minimum increment $\delta=0.01$. This ensures that if validation loss is not changing for more than 30 epochs by more than $\delta$, the training process stops. We utilize an ADAM optimizer with the same learning rate (0.001) for all three models.
@@ 4.4.3. Evaluation Metrics
The output of the model is the sleep label which is a binary variable for all the participants in the graph. We report the accuracy score (0-1), which is computed by dividing the total number of correct predictions by total predictions. To establish statistical significance between different model’s performance, we conduct an Analysis of Variance (ANOVA) test(Girden, 1992). ANOVA compares the performances of the four models across all trials and tests the null hypothesis that there is no statistically significant difference between the performance of different models. If the p-value is less than $\alpha=0.05$, we reject the null hypothesis. For further pair-wise comparison between proposed models and benchmarks, we conduct Tukey HSD (honestly significant difference) test(Tukey, 1977) for each pair between the proposed model and benchmarks.
5. Robustness Analysis against Data Perturbations
5.1. Artificial Perturbations
Figure 4. Different scenarios for perturbing the system. The circles represent nodes of the graph and the corresponding array represents a feature vector for that participant. The red arrow represents perturbation and red nodes indicate that those participants are perturbed. Yellow nodes are in their original state. (a) The original state of the system without any perturbation. (b) Randomly chosen 2 features $f_{2}$ and $f_{4}$ in red color are perturbed in all participants. (c) Randomly chosen features in further randomly chosen 2 participants $u_{2}$ and $u_{4}$ are perturbed.
The data are perturbed at two levels, feature level and participant level. For feature level perturbations, we assume that the feature $f_{i}$, where $i\in\{1,2,..m\}$, follows an underlying univariate Gaussian distribution, $\mathcal{N}(\mu_{f_{i}},\sigma_{f_{i}})$. After estimating the parameters of distribution from data, the perturbation is created by replacing feature data with randomly sampled values from the learned distribution with extended standard deviation. Note that if we create aggressive perturbations by replacing data with values that the feature does not often take, then those would be detected as outliers and filtered out in the preprocessing phase. Instead, we create noise and/or missingness by imputing values that would not be detected in the preprocessing phase or would be actually used for imputation if data are missing.
In the first scenario, we consider feature-level perturbation across all participants. First, a random subset $\tilde{I}_{f}$ of size $\tilde{m}<m$ is selected from feature index set $I_{f}=\{1,2,..,m\}$. Then, for each feature $f_{p}$, where $p\in\tilde{I}_{f}$, the values are imputed with randomly selected data from the corresponding distribution $\mathcal{N}(\mu_{f_{p}},\sigma_{f_{p}})$. This procedure is repeated for all participants in the network. This scenario is explained in Fig. 4(b). In this example, the first 2 features are chosen at random, $f_{2}$ and $f_{4}$. Then the features are perturbed in all participants as shown by red participant nodes.
5.2. Quantifying Robustness & Experiment Design
To characterize the impact of perturbations on model performance, we utilize the accuracy score as the metric. It is important to randomize and repeat the process multiple times to counter sensitivity to the choice of features and participants. First, the model is trained on unperturbed training data. Then, the perturbation is applied to test data and the accuracy score is computed for all test samples. The data are randomly perturbed again $H_{p}$ times and in each iteration, the performance metrics are computed. The original trained model is sensitive to parameter initialization and training-test split of data. To alleviate this sensitivity, we repeat the process of random splitting, training a new model and testing it on perturbed data $H_{t}$ times. Finally, for a given set of design parameters, the perturbation process is repeated $H_{t}\times H_{p}$ times. The average scores and $95\%$ confidence interval in these metrics for SMS graphs are reported in the results presented in the next subsection.
In order to further investigate how network topology and a participant’s location in the network impact its performance, we evaluate the importance of a node and its accuracy score. We characterize node importance by its centrality in the graph network(Easley and Kleinberg, 2010). To this end, we use two centrality metrics, degree, and eigenvalue centrality. Degree centrality assigns higher importance to nodes that have a large number of neighbors. However, it does not account for the cascade effect resulting from the fact that a node can also be important if it is connected to influential nodes. Eigenvalue centrality quantifies the influence of a node in a network by measuring the node’s closeness to influential parts of a network. More details on these metrics are provided in Supplementary Material B.1.
After computing absolute centrality metrics, we categorize them into three categories: alone, small, and large. Alone represents nodes that are not connected to anyone or have very small eigenvalue centrality ($C_{e}<10^{-4}$) and are mainly affected by perturbations in their own features. Small centrality nodes have small degree ($40<C_{d}<$ 4) or eigenvalue centrality ($0.310^{-4}<C_{e}<$ 0.3). Centrality scores greater than these thresholds are considered large. For each category, evaluation metrics are computed for a given trial and then average results over multiple ($\approx 500$) trials are reported.
@@ 6.1. Sleep duration prediction
Model Mean Accuracy Score (std)
Random Split LOCO
LSTM 0.75 (0.01) 0.71 (0.03)
CONV-LSTM 0.76 (0.05) 0.77 (0.02)
Proposed GCN-LSTM 0.81 (0.02) 0.83 (0.02)
Proposed GAN-LSTM 0.83 (0.02) 0.84 (0.02)
Table 2. Evaluation Results
ANOVA confirms that differences between models’ means for both types of the split are statistically significant (Random split p-value$=5.2x10^{-8}$, Leave One Cohort Out (LOCO) p-value$=6.3x10^{-15}$ )
The performance over a randomly chosen test set is evaluated in terms of accuracy score and the process is repeated 25 times for each graph size and temporal memory. A similar procedure is repeated for Leave-One-Cohort-Out (LOCO). The mean accuracy score and the standard deviation in accuracy across the trials are reported in Table 2. LSTM alone performs the worst among all models. Adding CNN slightly improves the performance because it uses the kernel to observe and slide across chunks of data from multiple participants. However, since there is no systematic mechanism for aggregating information from multiple participants, the improvement is very small. CONV-LSTM is also the most unstable among all given the high variation in the accuracy score. The proposed method GCN-LSTM incorporates additional information about inter-participant relationships and therefore the performance is improved significantly. Our final model, GAN-LSTM further incorporates attention while aggregating information from multiple participants resulting in the best performance. The variance in the performance of the proposed models is also small. ANOVA confirms that the difference in the performance of all four models is statistically significant. Further pair-wise comparison between proposed models and benchmarks through Tukey HSD results in all adjusted p-values$<\alpha=0.05$. This establishes that the proposed models perform statistically significantly better than the benchmarks. The exact p-values are reported in Supplementary material Table. 3.
@@ 7.1. Perturbations in a subset of features
In the first experiment, we perturb randomly selected features in inputs for all participants in a graph of size 15 and calculate performance. By increasing the percentage of perturbed features, we average the results over 500 trials, as shown in Fig. 6(a). At x=0 (no perturbation), all models perform well. As perturbations increase, performance drops, with LSTM showing overall poor performance. CONV-LSTM exhibits a sharp performance decline and is the least robust with numerous perturbed features. GCN also shows decreasing performance with more disturbance. However, attention-based GAN remains highly robust against perturbations.
@@ 7.2. Robustness and network topology
In the second experiment, instead of all participants experiencing the same perturbation, a random subset of participants in a network is chosen. This scenario is depicted in Fig. 4 (c). This setting, where only some participants of the network experience disturbance, is more frequent in real-life data. While aggregating information from multiple participants has shown promising improvement in prediction accuracy, it also makes the predictions of the entire network more vulnerable to perturbations if a few participants in the network have noisy/missing data.
Thus, we present how the proportion of perturbed participants impacts the performance in Fig. 6(b) by perturbing $50\%$ features in a varying number of participants. As observed in Fig 6(b) LSTM experiences very high errors consistently. CONV-LSTM shows a steep increase in error as more and more participants are perturbed. Accuracy score as a metric is itself slightly robust to small changes in predictions that do not alter the prediction category. To identify the absolute behavior of the model, we also report the error metric MAPE in Supplementary Material Figure 11. In Fig. 11(b), when 13 out of 15 participants are perturbed, there is a sharp drop in the performance of the CONV-LSTM model. GCN also experiences a steeper increase in error. However, the attention-based model is still quite robust to these perturbations.
@@
The results presented in Figures 6 represent an average impact on the entire network. However, network topology also plays a role in how the model’s performance for some participants is more robust to noisy or missing feature data than others. In order to segregate these different parts of the network, we report the average accuracy score of different degree category nodes when 5 out of 15 participants experience perturbations in $50\%$ features, both chosen randomly, in Fig.7(a). The average performance without any noise for each model is represented by dashed lines. The category ’alone’ indicates nodes that are not connected to anyone. The drop in performance in the ’alone’ category results from the vulnerability of a participant to noise in its own features. Both GCN and CONV-LSTM suffer from a significant drop in performance for alone nodes. For nodes with a small degree, CONV-LSTM experiences a drop while other models are robust to it. Finally, for participants with a large number of direct neighbors, all models observe a drop in accuracy score because they have a higher chance of being closely connected to a node that is perturbed. We further increase the perturbations in Fig.7 by perturbing 13 out of 15 participants. The drop in performance is higher than what is observed in Fig.7(a), especially for large degree nodes.
@@
Figure 7. Impact on the performance of different nodes in the network by perturbing $50\%$ features in graphs of size 15 and sequence length of 3 days. The dashed lines represent the average accuracy score without any perturbation.
@@
Figure 8. Impact on the performance of different nodes in the network by perturbing $50\%$ features in graphs of size 15 and sequence length of 3 days. The dashed lines represent the average accuracy score without any perturbation.
@@
We use another centrality metric to segregate different parts of the network known as eigenvalue centrality in Fig.8. In the first experiment, a small number of participants i.e., 5 are perturbed in Fig.8(a). For participants who are alone or have extremely small eigenvalue centrality ($C_{e}<10^{-4}$), there is a sharp drop in performance. For small eigenvalues, only CONV-LSTM suffers from a drop in the accuracy score. For large eigenvalue, the behavior is unexpected and instead of a decrease, there is an increase in average performance. To investigate it further, we raised the number of perturbed participants to 13 (Fig.8(b)). For the ’alone’ category, the drop in performance is slightly larger than before. For small eigenvalue centrality, we observe unexpected behavior, where performance is slightly improving for graph-based models. Finally, for large centrality nodes, there is a significant decrease in performance. This category is the only one in which even attention-based models are not robust. These results indicate that more than connections to influential nodes can cause unstable behavior in model predictions.
@@
Finally, we investigate how perturbations in different parts of the sequence impact the performance. For CONV-LSTM and GCN, as the proportion of perturbed days increases, the performance drops in a linear fashion. Attention-based models are robust to these perturbations. Detailed results and discussion are provided in Supplementary Material section C.3.1.

### 27 Jan 2024 v2

@@ 4.1. Problem Description
We pose a problem of predicting sleep labels for the next day based on wearable and mobile data for the current and past few days. Formally, the multi-modal data with $m$ features for a given day $k$ is represented by the vector $x[k]\in\mathbb{R}^{m}$ and the corresponding sleep duration in minutes is represented by $s[k]$. The feature matrix $X[k]$ is composed of past $L$ days of feature data, where $L$ is an indicator of temporal memory, and the sleep label $y[k]$ is generated from sleep duration $s[k]$ (mins). The recommended sleep duration for young adults is between $8-10$ hours (Paruthi et al., 2016; Hirshkowitz et al., 2015b, a), so we choose a threshold of 8 hours for differentiating between good and poor sleep behavior. Also, the sleep duration distribution after preprocessing is relatively balanced around 8 hours. The sleep duration for our collected dataset had a mean of 7.2 hours (std= 2.1 hours).
@@ 4.2. Graph Network Construction
We validated social interactions among study participants using data from pre- and post-study social network surveys. These data allowed us to quantify the closeness between participants and create graph networks, which are detailed in Supplementary Material section E.2. Our findings affirm the presence of social interactions and phone communication among the study participants.
@@
4.3. SleepNet: Proposed Architecture
@@ 4.3.2. Attention-based Context Aggregation
To integrate graph structure in this attention mechanism, these coefficients are only computed for the first-hop neighbors. In addition, to compare different neighbors and have stable information aggregation, the coefficients are normalized across all neighbors. Finally, the normalized coefficients are used to compute a weighted aggregation of features from all over the node $u_{i}$’s neighbors.
@@ 4.3.5. End-to-End Model
Figure 3. SleepNet Architecture. The top module based on GAN takes both the feature matrix and graph as input. The layer operates on graphs by applying attention shown by blue dashed links and original graph links are shown in solid green lines. The bottom module takes sequential data as input and extracts temporal trends. Finally, embedded features from both modules are combined in the aggregation module that produces predicted sleep labels.
The complete architecture of SleepNet is shown in Fig.3. The input to the model is a 3-tuple object $I=(X[k],S^{L}_{k},A_{N}^{k})$, where $X[k]=\left[x_{1}[k],x_{2}[k],...x_{N}[k]\right]$ is a matrix containing feature vectors of $N$ participants each represented by $x_{i}\in\mathbb{R}^{m}$ for day indexed by $k$. The sequence data fed to the LSTM unit is represented by $S^{L}_{k}$. The graph network corresponding to the $N$ participants in feature matrix $X[k]$ is represented by adjacency matrix $A_{N}^{k}$. This corresponding data label is sleep duration for all $N$ participants for the next day represented by $y[k+1]$. After extracting temporal dynamics and multi-participant aggregated features in an embedded space, the output of both modules is normalized and concatenated before being fed to a small network of dense layers that produce the final output.
@@ 4.4.2. Machine Learning Pipeline
  • Random Split: The dataset is randomly split into training and test sets. We use random splitting to avoid selection bias. Around $50\%$ of data is used for training, $10\%$ for validation, and the remaining $40\%$ for testing with SMS graphs. The same training data is used to train all 4 models and then the same test data is used for testing across all models. This ensures that a fair comparison is conducted between the models. In order to mitigate the impact of model sensitivity to parameter initialization and depict the stability of the model, we use bootstrapping (Efron, 1992) and repeat the training and testing procedure 25 times and report average results along with the standard deviation across trials.
@@
To reduce the computation time, we use early stopping with a patience value of 30 epochs and minimum increment $\delta=0.01$. This ensures that if validation loss is not changing for more than 30 epochs by more than $\delta$, the training process stops. We utilize an ADAM optimizer with the same learning rate (0.001) for all three models.
@@ 4.4.3. Evaluation Metrics
The output of the model is the sleep label which is a binary variable for all the participants in the graph. We report the accuracy score (0-1), which is computed by dividing the total number of correct predictions by total predictions. To establish statistical significance between different model’s performance, we conduct an Analysis of Variance (ANOVA) test (Girden, 1992) and Kruskal–Wallis non parametric test(McKight and Najab, 2010). These tests compare the performances of the four models across all trials and test the null hypothesis that there is no statistically significant difference between the performance of different models. If the p-value is less than $\alpha=0.05$, we reject the null hypothesis. For further pair-wise comparison between proposed models and benchmarks, we conduct Tukey HSD (honestly significant difference) test (Tukey, 1977) for each pair between the proposed model and benchmarks.
@@ 5.1. Sleep Duration Prediction
Table 1. Evaluation Results.
Model Mean Accuracy Score (std) Random Split LOCO LSTM 0.75 (0.01) 0.71 (0.03) CONV-LSTM 0.76 (0.05) 0.77 (0.02) Proposed GCN-LSTM 0.81 (0.02) 0.83 (0.02) Proposed GAN-LSTM 0.83 (0.02) 0.84 (0.02)
Model Benchmark
p-value
(Random
Split)
p-value
GCN-LSTM LSTM $1.0x10^{-4}$ 0.00
GCN-LSTM CONV-LSTM $2.0x10^{-4}$ $1.0x10^{-3}$
GAN-LSTM LSTM 0.00 0.00
GAN-LSTM CONV-LSTM 0.00 0.00
Table 1. Evaluation Results.
Table 2. Tukey-HSD adjusted p-values.
The performance over a randomly chosen test set is evaluated in terms of accuracy score and the process is repeated 25 times for each graph size and temporal memory. A similar procedure is repeated for Leave-One-Cohort-Out (LOCO). The mean accuracy score and the standard deviation in accuracy across the trials are reported in Table 2. ANOVA reveals that differences between models’ means for both types of the split are statistically significant (Random split p-value$=5.2x10^{-8}$, LOCO p-value$=6.3x10^{-15}$). Additionally, the Kruskal-Wallis non-parametric test also confirms that there is a significant difference among different models (Random split p-value$=3.2x10^{-6}$, LOCO p-value$=4.5x10^{-9}$).
LSTM alone performs the worst among all models. Adding CNN slightly improves the performance because it uses the kernel to observe and slide across chunks of data from multiple participants. However, since there is no systematic mechanism for aggregating information from multiple participants, the improvement is very small. CONV-LSTM is also the most unstable among all given the high variation in the accuracy score. The proposed method GCN-LSTM incorporates additional information about inter-participant relationships and therefore the performance is improved significantly. Our final model, GAN-LSTM further incorporates attention while aggregating information from multiple participants resulting in the best performance. The variance in the performance of the proposed models is also small. ANOVA confirms that the difference in the performance of all four models is statistically significant. Further pair-wise comparison between proposed models and benchmarks through Tukey HSD results in all adjusted p-values$<\alpha=0.05$ as shown in Table 2. This establishes that the proposed models perform statistically significantly better than the benchmarks.
@@ 5.2. Impact of Network Characteristics
5.3. Feature Importance
Figure 5. Feature Importance for different modalities for N=10 and L=3 days.
SC, SCL, and SD stand for Skin Conductance, Skin Conductance Level, and Standard Deviation respectively.
More details on features, in Tables 3,4, and 5.
We further investigate the impact of the temporal module by removing it and observing the drop in performance. Our findings highlight that the temporal module aids in improving performance in proposed architectures and its impact is more pronounced in GCN-LSTM than GAN-LSTM. More details in Supplementary Material section D.
6. Robustness Analysis against Data Perturbations
6.1. Artificial Perturbations
Figure 6. Different scenarios system perturbation. The circles represent nodes/users in the graph and the corresponding array represents the feature vector. The red arrows and red nodes represents perturbed participants. Yellow nodes have no perturbation. (a) The original system without any perturbation. (b) Randomly chosen 2 features $f_{2}$ and $f_{4}$ in red color are perturbed in all participants. (c) Randomly chosen features in 2 randomly chosen participants $u_{2}$ and $u_{4}$ are perturbed.
The data are perturbed at two levels, feature level and participant level. For feature level, we assume that the feature $f_{i}$, where $i\in\{1,2,..m\}$, follows an univariate Gaussian distribution, $\mathcal{N}(\mu_{f_{i}},\sigma_{f_{i}})$. After estimating the distribution parameters from data, perturbations are created by replacing feature data with randomly sampled values from the learned distribution with extended standard deviation. Note that if we create aggressive perturbations by replacing data with values that the feature does not often take, then those would be detected as outliers and filtered out in the preprocessing phase. Instead, we create noise and/or missingness by imputing values that would not be detected in the preprocessing phase or would be actually used for imputation if data are missing.
In the first scenario shown in Fig. 6(b), we consider feature-level perturbation across all participants. First, a random subset $\tilde{I}_{f}$ of size $\tilde{m}<m$ is selected from feature index set $I_{f}=\{1,2,..,m\}$. Then, for each feature $f_{p}$, where $p\in\tilde{I}_{f}$, the values are imputed with randomly selected data from the corresponding distribution $\mathcal{N}(\mu_{f_{p}},\sigma_{f_{p}})$. This procedure is repeated for all participants in the network. In this example, the first 2 features are chosen at random, $f_{2}$ and $f_{4}$. Then the features are perturbed in all participants as shown by red participant nodes.
6.2. Quantifying Robustness & Experiment Design
To characterize the impact of perturbations on model performance, we utilize the accuracy score as the metric. It is important to randomize and repeat the process multiple times to counter sensitivity to the choice of features and participants. First, the model is trained on unperturbed training data. Then, the perturbation is applied to test data and the accuracy score is computed for all test samples. The data are randomly perturbed again $H_{p}$ times and in each iteration, the performance metrics are computed. The original trained model is sensitive to parameter initialization and training-test split of data. To alleviate this sensitivity, we repeat the process of random splitting, training a new model and testing it on perturbed data $H_{t}$ times. Finally, for a given set of design parameters, the perturbation process is repeated $H_{t}\times H_{p}$ times. The average scores and $95\%$ confidence interval are reported.
To further investigate how network topology and a participant’s location in the network impact its performance, we evaluate the importance of a node and its accuracy score. We characterize node importance by its centrality in the graph network (Arney, 2013). To this end, we use two centrality metrics, degree, and eigenvalue centrality. Degree centrality assigns higher importance to nodes that have a large number of neighbors. However, it does not account for the cascade effect resulting from the fact that a node can also be important if it is connected to influential nodes. Eigenvalue centrality quantifies the influence of a node in a network by measuring the node’s closeness to influential parts of a network. More details on these metrics are provided in Supplementary Material B.1.
The centrality metrics are categorized into three categories: alone, small, and large. Alone nodes are not connected to anyone or have very small eigenvalue centrality ($C_{e}<10^{-4}$) and are mainly affected by perturbations in their own features. Small centrality nodes have small degree ($40<C_{d}<$ 4) or eigenvalue centrality ($0.310^{-4}<C_{e}<$ 0.3). Centrality scores greater than these thresholds are considered large. For each category, evaluation metrics are computed for a given trial and then average results over multiple ($\approx 500$) trials are reported.
@@ 7.1. Perturbations in a subset of features
In the first experiment, we perturb randomly selected features in inputs for all participants in a graph of size 15 and calculate performance. By increasing the percentage of perturbed features, we average the results over 500 trials, as shown in Fig. 7(a). Features were perturbed in increments of 25 (e.g., 100, 125, 150, …), and then converted to percentages for better understanding. At x=0 (no perturbation), all models perform well. As perturbations increase, performance drops, with LSTM showing overall poor performance. CONV-LSTM exhibits a sharp performance decline and is the least robust with numerous perturbed features. GCN also shows decreasing performance with more disturbance. However, attention-based GAN remains highly robust against perturbations.
@@ 7.2. Robustness and Network Topology
In the second experiment depicted in Fig. 6 (c), instead of all participants experiencing the same perturbation, a random subset of participants in a network is chosen. While aggregating information from multiple participants has shown promising improvement in prediction accuracy, it also makes the predictions of the entire network more vulnerable to perturbations if a few participants in the network have noisy/missing data. Thus, we present how the proportion of perturbed participants impacts the performance in Fig. 7(b) by perturbing $50\%$ of the features in a varying number of participants. As observed in Fig 7(b) LSTM experiences very high errors consistently. CONV-LSTM shows a steep increase in error as more and more participants are perturbed.
@@
The results presented in Figures 7 represent an average impact on the entire network. However, network topology also plays a role in how the model for some participants is more robust to noisy or missing feature data than others. In order to segregate these different parts of the network, we report the average accuracy score of different degree category nodes when 5 out of 15 participants experience perturbations in $50\%$ features, both chosen randomly, in Fig.8(a). The average performance without any noise for each model is represented by dashed lines. The category ’alone’ indicates nodes that are not connected to anyone. The drop in performance in the ’alone’ category results from the vulnerability of a participant to noise in its own features. Both GCN and CONV-LSTM suffer from a significant drop in performance for alone nodes. For nodes with a small degree, CONV-LSTM experiences a drop while other models are robust to it. Finally, for participants with a large number of direct neighbors, all models observe a drop in accuracy score because they have a higher chance of being closely connected to a node that is perturbed. We further increase the perturbations in Fig.8 to 13 participants. The drop in performance is higher than what is observed in Fig.8(a), especially for large degree nodes.
@@
Figure 8. Impact on the performance of different degree centrality nodes in the network by perturbing $50\%$ features in graphs of size 15 and sequence length of 3 days. The dashed lines represent the average accuracy score without any perturbation.
@@
Figure 9. Impact on the performance of different eigenvalue centrality nodes by perturbing $50\%$ features in graphs of size 15 and sequence length of 3 days. The dashed lines represent the average accuracy score without any perturbation.
@@
We use another centrality metric called eigenvalue centrality to segregate different parts of the network. First, a small number of participants i.e., 5 are perturbed in Fig.9(a). For participants who are alone or have extremely small eigenvalue centrality ($C_{e}<10^{-4}$), there is a sharp drop in performance. For small eigenvalues, only CONV-LSTM suffers from a drop in the accuracy score. For large eigenvalue, the behavior is unexpected and instead of a decrease, there is an increase in average performance. To investigate it further, we raised the number of perturbed participants to 13 (Fig.9(b)). For the ’alone’ category, the drop in performance is slightly larger than before. For small eigenvalue centrality, we observe unexpected behavior, where performance is slightly improving for graph-based models. Finally, for large centrality nodes, there is a significant decrease in performance. This category is the only one in which even attention-based models are not robust. These results indicate that more connections to influential nodes can lead to unstable predictions.
@@
Finally, we investigate how perturbations in different parts of the sequence impact the performance. Performance declines linearly for CONV-LSTM and GCN with increasing sequence perturbations, while attention-based models demonstrate robustness to these changes.Detailed results are provided in Supplementary section C.2.1.